The length of the top of the table is 6m greater than the width. The area is 72m squared. Find the dimensions of the table
Answer by the_tutor2011(8)

x(x+6)=72
x^2+6x=72
x^2+6x-72=0
(x+12)(x-6)=0
x=-12 or x=6
width = 6
length = 12